solve for x in the two equations below using substitution
12x - 3y = 36
24x + 3y= 180

Answer by richwmiller(17219)

Substitution is the wrong method to use.
Addition or elimination is easier.
12x - 3y = 36
24x + 3y= 180
add
36x=216
x=6
substitution
12x - 3y = 36
12x-36=3y
24x + 3y= 180
3y=180-24x
12x-36=180-24x
36x-36=180
36(x-1)=180
x-1=5
x=6